# ProtoSiTex: Learning Semi-Interpretable Prototypes for Multi-label Text Classification

### v1

@@ Abstract.
The surge in user-generated reviews has amplified the need for interpretable models that can provide fine-grained insights. Existing prototype-based models offer intuitive explanations but typically operate at coarse granularity (sentence or document level) and fail to address the multi-label nature of real-world text classification. We propose ProtoSiTex, a semi-interpretable framework designed for fine-grained multi-label text classification. ProtoSiTex employs a dual-phase alternating training strategy: an unsupervised prototype discovery phase that learns semantically coherent and diverse prototypes, and a supervised classification phase that maps these prototypes to class labels. A hierarchical loss function enforces consistency across sub-sentence, sentence, and document levels, enhancing interpretability and alignment. Unlike prior approaches, ProtoSiTex captures overlapping and conflicting semantics using adaptive prototypes and multi-head attention. We also introduce a benchmark dataset of hotel reviews annotated at the sub-sentence level with multiple labels. Experiments on this dataset and two public benchmarks (binary and multi-class) show that ProtoSiTex achieves state-of-the-art performance while delivering faithful, human-aligned explanations, establishing it as a robust solution for semi-interpretable multi-label text classification.
Text Classification, Prototype Learning, Semi-interpretable Models
journalyear: 2025
1. Introduction
@@
In recent years, online user reviews have become influential across domains like e-commerce, hospitality, and social media (Diaz et al., 2018). These reviews offer rich, unstructured insights into user preferences (Duan et al., 2008; Banerjee et al., 2017; Floyd et al., 2014). Extracting meaningful information from them is critical, and transformer-based language models (LMs) such as BERT (Devlin et al., 2019) have excelled in NLP tasks like sentiment analysis (Wankhade et al., 2022), question answering (Rogers et al., 2023), and text classification (Minaee et al., 2021). However, these models are typically opaque, limiting interpretability, an essential factor for trust, fairness, and compliance (Rudin, 2019; Castelvecchi, 2016). To address the limitations of black-box models, post-hoc explanation methods such as LIME (Ribeiro et al., 2016) and SHAP (Lundberg et al., 2017) have been proposed. However, they often lack faithfulness and consistency, offering only approximate and sometimes inaccessible justifications (Crabbé et al., 2023; Alvarez-Melis et al., 2018), especially for non-technical users. This has led to growing interest in interpretable models that provide transparency by design (Molnar, 2020).
@@
Prototype-Based Networks (PBNs) follow the paradigm of inherent interpretability, mimicking human reasoning via similarity to class-representative prototypes (Rosch, 1975). While effective in vision (Chen et al., 2019; Nauta et al., 2023) and recently adapted to NLP (Hong et al., 2023; Wen et al., 2025), most PBNs are limited to single-label, sentence- or document-level tasks. This limits their use in real-world, multi-label classification scenarios, e.g., “The view was breathtaking, but the staff was rude” expresses opposing sentiments, challenging coarse-grained prototype models (Sourati et al., 2024).
@@
Efforts to improve interpretability in text classification often come at the expense of predictive strength. Transformer-based models capture rich semantic structures and typically outperform simpler methods in benchmarks (Li et al., 2022), yet their lack of transparency limits trust in sensitive applications. Prototype-driven approaches (Wei et al., 2024; Wen et al., 2025) provide clearer, case-based explanations that are easier for users to follow. However, they may sacrifice raw accuracy or struggle to scale to complex domains. This creates a persistent issue: optimizing solely for accuracy risks overlooking fairness and accountability, while prioritizing interpretability may reduce competitiveness in performance-driven evaluations. The issue becomes even more pronounced in multi-label classification (Pontiki et al., 2016; Schouten et al., 2015), where models must balance fine-grained predictive capability with explanations that remain faithful and comprehensible.
@@
To address these limitations, we propose ProtoSiTex, a semi-interpretable prototype-based framework for fine-grained multi-label classification. ProtoSiTex effectively handles overlapping sentiments and multi-label scenarios via three key innovations: (i) adaptive prototype learning using multi-head attention for sub-sentence-level semantic alignment, (ii) a dual-phase alternating training strategy that combines unsupervised prototype discovery with supervised classification to improve generalization and reduce overfitting, and (iii) a hierarchical loss function that enforces consistency across sub-sentence, sentence, and document levels for structured and faithful predictions. We use the term semi-interpretable to denote models that provide faithful, prototype-based explanations for their predictions while still relying on latent neural mechanisms (e.g., attention, embeddings, hierarchical aggregation) that are not inherently transparent. Our key contributions are:
@@
(i) A semi-interpretable classification framework using adaptive prototypes and multi-head attention: We introduce ProtoSiTex, a prototype-based model primarily for multi-label classification with interpretability. Unlike prior PBNs limited to sentence or document granularity, ProtoSiTex learns adaptive prototypes aligned with sub-sentence-level class patterns, enabling it to disentangle overlapping or conflicting sentiments. Leveraging multi-head attention, it captures diverse semantic interactions between text segments and prototypes, enhancing both alignment and interpretability.
@@
(ii) Dual-phase alternating training strategy and hierarchical loss for robust learning and semantic consistency: To ensure generalizable and interpretable prototypes, we propose a dual-phase alternating learning strategy: (a) an unsupervised prototype discovery phase that promotes diversity and semantic coherence via clustering, and (b) a supervised classification phase that maps prototypes to multiple labels. This decoupling reduces overfitting by allowing periodic prototype refinement. We also introduce a hierarchical loss function to enforce consistency across sub-sentence, sentence, and document levels, enhancing structured reasoning, predictive accuracy, and explanation fidelity.
@@
(iii) New dataset and extensive evaluation: We curate a hotel review dataset (HR) annotated at the sub-sentence-level with multi-label aspect tags, addressing the limitations of existing sentence or document-label benchmarks. To evaluate ProtoSiTex, we conducted extensive experiments on three datasets: (a) HR (multi-label classification), (b) IMDb (binary classification), and (c) TweetEval (multi-class classification). Across all datasets, ProtoSiTex consistently outperformed existing interpretable baselines and achieved performance on par with state-of-the-art black-box models, showing its effectiveness in terms of accuracy, robustness, and interoperability.
2. Related Work
@@
Interpretable Text Classification: Deep models like CNNs, RNNs, and transformers have achieved strong performance across NLP tasks (Li et al., 2022), but their opaque nature hinders transparency. Post-hoc explanation methods, such as LIME (Ribeiro et al., 2016) and SHAP (Lundberg et al., 2017), aim to interpret predictions by approximating local feature importance. However, these techniques operate independently of model training and often produce inconsistent or non-faithful rationales (Cesarini et al., 2024). This has spurred interest in interpretable models that learn to classify and explain simultaneously.
@@
Prototype-based Networks: Prototype-Based Networks (PBNs) classify instances based on similarity to learned prototype vectors, offering human-aligned explanations via exemplar-based reasoning (Snell et al., 2017). Early works such as ProSeNet (Ming et al., 2019) and ClassVector (Yao et al., 2019a) introduced prototype learning for document-level, single-label tasks with limited interpretability. Subsequent methods like ProtoryNet (Hong et al., 2023) and ProtoLens (Wei et al., 2024) learned finer-grained prototypes at the sentence or subsentence level. GAProtoNet (Wen et al., 2025) integrated graph-attention with prototypes for relational reasoning. Yet, most PBNs focus on single-label classification, lacking the granularity and compositionality required for multi-label classification.
@@
Multi-Label Text Classification: Multi-label text classification (MLTC) involves assigning multiple labels to a document, making it more challenging than single-label tasks due to inherent label dependencies and data imbalance (Liu et al., 2021). Early approaches relied primarily on problem transformation methods, such as Binary Relevance, Label power set, and Classifier Chains, as well as algorithm adaptation methods, including ML-SVMs and ML-kNN (Zhang et al., 2013). Models based on CNN and RNN improved text representation learning (Yang et al., 2016; Liu et al., 2017), while Graph Convolutional Networks were introduced to explicitly capture label correlations (Ma et al., 2021). More recently, Transformer-based architectures, including BERT (Devlin et al., 2019) have emerged as strong baselines. Despite these advances, existing methods either treat MLTC as a flat classification problem or rely on complex black-box architectures with limited interpretability. Furthermore, few datasets provide sub-sentence-level multi-label annotations, limiting empirical benchmarking.
@@
Positioning: ProtoSiTex overcomes these limitations by unifying prototype-based interpretability with fine-grained multi-label classification. Unlike prior works, it learns adaptive prototypes aligned with sub-sentence semantics and leverages a dual-phase learning strategy, combining unsupervised prototype discovery and supervised hierarchical classification. This design enables structured predictions from sub-sentence to document level while providing reliable, human-understandable explanations.
3. Proposed Methodology
@@
Problem formulation: Given a document $\mathcal{D}$ composed of multiple sentences, our goal is to predict its associated class(es), accommodating both single-label and multi-label classification. ProtoSiTex performs document-level prediction while highlighting the specific content segments responsible for each label.
@@
Preprocessing: In our framework, sub-sentence annotations, though optional, enhance fine-grained interpretability in multi-label tasks. Since manual labeling is costly, we use a few-shot LLM-based annotation engine (Dutta et al., 2025) to infer them from document-level labels automatically.
@@
Each document $\mathcal{D}$ comprises $n$ sentences, represented as $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots,\mathcal{S}_{n}\}$. Each sentence $\mathcal{S}_{i}$ is further segmented into $t_{i}$ sub-sentences, defined as $\mathcal{S}_{i}=\{s_{i1},s_{i2},\ldots,s_{it_{i}}\}$, where punctuation marks such as commas and semicolons are used as segmentation cues. The total number of sub-sentences in a document is denoted by $t=\sum_{i=1}^{n}t_{i}$. To assign fine-grained class labels, we apply the LLM-based engine to generate multi-label annotations at the sub-sentence level. These annotations are encoded in a binary label matrix $\mathcal{C}\in\{0,1\}^{t\times l}$, where $l$ is the total number of classes. A matrix entry $c_{ij}=1$ indicates that the $i^{\text{th}}$ sub-sentence is associated with the $j^{\text{th}}$ label. For hierarchical supervision, we aggregate sub-sentence labels to form sentence-level labels, while document-level labels are directly provided. This bottom-up strategy ensures consistent multi-level annotation. To support structured supervision and hierarchical loss modeling, we construct two binary mapping matrices. The first, $\mathcal{M}^{1}\in\{0,1\}^{1\times n}$, captures the mapping from a document to its sentences. The second, $\mathcal{M}^{2}\in\{0,1\}^{n\times t}$, captures the mapping from each sentence to its associated sub-sentences. An entry $\mathcal{M}^{*}_{ij}=1$ indicates that sentence or sub-sentence $j$ belongs to sentence or document $i$, respectively.
Figure 1. Proposed solution architecture: ProtoSiTex
3.1. ProtoSiTex Architecture
@@
The proposed framework, ProtoSiTex (Fig. 1), consists of three core stages: (i) encoding sub-sentences into high-dimensional semantic representations using a language model (LM)-based encoder, (ii) employing a dual-phase learning strategy that first discovers representative prototypes through unsupervised learning, then performs supervised classification guided by hierarchical supervision using document-level taxonomy, and (iii) performing inference through prototype-based reasoning applied to the learned embeddings. We detail each component below.
@@
Sub-sentence Representation via LM-based Encoder: Effective text classification relies on context-aware representations. While traditional embeddings (e.g., Word2Vec, GloVe, FastText) are static, LM-based encoders offer dynamic, context-sensitive embeddings but lack interpretability. To balance expressiveness and transparency, we use an LM-based encoder to generate sub-sentence embeddings.
@@
For each $s_{ij}\in\mathcal{S}_{i}$, the encoder outputs $e_{ij}\in\mathbb{R}^{d}$, producing a feature matrix $\mathcal{E}\in\mathbb{R}^{t\times d}$ over $t$ sub-sentences. This matrix feeds into the prototype discovery and classification modules, enabling both semantic precision and interpretability. We next describe the dual-phase learning.
@@
Clustering-phase for Prototypes Learning: The first stage of our dual-phase learning aims to discover semantically meaningful and interpretable prototypes that facilitate multi-label classification. To achieve this, we introduce a learnable prototype layer coupled with a multi-head attention (MHA) mechanism, enabling diverse alignment between sub-sentence embeddings and prototypes.
@@
We initialize $q$ prototypes $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{q}\};~\forall p_{i}\in\mathbb{R}^{d}$ using random initialization, which empirically outperforms K-Means, Xavier, and Kaiming. The value of $q$ is selected via grid search. Sub-sentence embeddings $\mathcal{E}\in\mathbb{R}^{t\times d}$ are projected into the prototype space using MHA, which learns flexible, context-aware alignments. For each attention head $i$, we compute query $\mathcal{Q}_{i}=\mathcal{P}\cdot\mathcal{W}^{{Q}_{i}}$, key $\mathcal{K}_{i}=\mathcal{E}\cdot\mathcal{W}^{{K}_{i}}$, and value $\mathcal{V}_{i}=\mathcal{E}\cdot\mathcal{W}^{{V}_{i}}$ matrices, where $\mathcal{W}^{Q_{i}},\mathcal{W}^{K_{i}},\mathcal{W}^{V_{i}}\in\mathbb{R}^{d\times d}$ are learnable parameters. Scaled dot-product attention computes the alignment:
@@
(1) $\mathcal{H}_{i}=\sigma\left(\left(\mathcal{Q}_{i}\cdot\mathcal{K}_{i}^{\top}\right)/\sqrt{d}\right)\cdot\mathcal{V}_{i};\hskip 16.38895pt\mathcal{H}=\left(\mathcal{H}_{1}~\lVert~\mathcal{H}_{2}~\lVert...\lVert~\mathcal{H}_{h}\right)\cdot\mathcal{W}^{h}$
@@
where, $\sigma$ denotes softmax function. The resulting matrix $\mathcal{H}\in\mathbb{R}^{q\times d}$ encodes prototype-aware representations of sub-sentences, which are used in the subsequent classification phase to associate prototypes with class labels.
@@
Classification-phase for Learning Multi-granular Document Details: In the second phase, we train the classifier using prototype embeddings $\mathcal{P}$, sub-sentence embeddings $\mathcal{E}$, and prototype-aware representations $\mathcal{H}$. To enhance generalization, we apply multi-granular supervision without separate classifiers, enabling two-way explainability: from sub-sentences to prototypes and prototypes to labels. To map prototypes to class labels, we pass $\mathcal{H}$ through two dense layers with ReLU followed by a softmax function, producing prediction scores $\tilde{\mathcal{G}}_{1}\in\mathbb{R}^{q\times l}$:
@@
(2) $\scriptsize\tilde{\mathcal{G}}_{1}={\sigma}\left(\mathcal{W}_{3}\cdot ReLU\left(\mathcal{W}_{2}\cdot ReLU\left(\mathcal{W}_{1}\cdot\mathcal{H}\right)\right)\right)$
@@
Using $\tilde{\mathcal{G}}_{1}$, we derive sentence-level predictions $\tilde{\mathcal{G}}_{2}=\text{Clamp}\left(\mathcal{M}^{2}\cdot\mathcal{Y}^{\top}\cdot\tilde{\mathcal{G}}_{1}\right)$ $\in\mathbb{R}^{n\times l}$ by aggregating sub-sentence contributions. This is done via the binary prototype-to-subsentence assignment matrix $\mathcal{Y}$, computed by applying cosine similarity between prototypes and sub-sentences, and sentence-to-subsentence mapping $\mathcal{M}^{2}$. Finally, document-level predictions $\tilde{\mathcal{G}}_{3}=\text{Clamp}\left(\mathcal{M}^{1}\cdot\tilde{\mathcal{G}}_{2}\right)\in\mathbb{R}^{1\times l}$ are derived by aggregating sentence-level scores via document-sentence mapping $\mathcal{M}^{1}$. The $\text{Clamp}(\cdot)$ function restricts the elements of a matrix to the $[0,1]$ range by applying element-wise clipping, setting any negative value to 0 and any value above 1 to 1.
@@
We generate dynamic ground-truths for training. Prototype-label associations $\mathcal{G}_{1}$ are computed by first measuring cosine similarity between prototypes and sub-sentence embeddings to form $\mathcal{X}=\mathcal{P}\cdot\mathcal{E}^{\top}/(\lVert\mathcal{P}\rVert\lVert\mathcal{E}^{\top}\rVert)$. This is binarized into $\mathcal{Y}\in\{0,1\}^{q\times t}$ as:
@@
(3) $\scriptsize\mathcal{Y}_{ij}=\begin{cases}1,&\text{if }i=\arg\max_{k\in\{1~,\dots,~q\}}\mathcal{X}_{kj},\\ 0,&\text{otherwise}\end{cases}$
@@
This binarized $\mathcal{Y}$ then multiplied with the sub-sentence label matrix $\mathcal{C}\in\{0,1\}^{t\times l}$ to yield:
(4) $\scriptsize\mathcal{G}_{1}=\mathcal{Y}\cdot\mathcal{C};\qquad\mathcal{G}_{2}=\text{Clamp}(\mathcal{M}^{2}\cdot\mathcal{C})$

### v2

@@ Abstract
The surge in user-generated reviews has amplified the need for interpretable models that can provide fine-grained insights. Existing prototype-based models offer intuitive explanations but typically operate at coarse granularity (sentence or document level) and fail to address the multi-label nature of real-world text classification. We propose ProtoSiTex, a semi-interpretable framework designed for fine-grained multi-label text classification. ProtoSiTex employs a dual-phase alternate training strategy: an unsupervised prototype discovery phase that learns semantically coherent and diverse prototypes, and a supervised classification phase that maps these prototypes to class labels. A hierarchical loss function enforces consistency across subsentence, sentence, and document levels, enhancing interpretability and alignment. Unlike prior approaches, ProtoSiTex captures overlapping and conflicting semantics using adaptive prototypes and multi-head attention. We also introduce a benchmark dataset of hotel reviews annotated at the subsentence level with multiple labels. Experiments on this dataset and two public benchmarks (binary and multi-class) show that ProtoSiTex achieves state-of-the-art performance while delivering faithful, human-aligned explanations, establishing it as a robust solution for semi-interpretable multi-label text classification.
I Introduction
@@
In recent years, online user reviews have become influential across domains like e-commerce, hospitality, and social media [review_helpfulness]. These reviews offer rich, unstructured insights into user preferences [online_review, online_review_1, online_review_2]. Extracting meaningful information from reviews is critical, and transformer-based language models (LMs) [devlin-etal-2019-bert] have excelled in NLP tasks like sentiment analysis [sentiment_survey], question answering [QA_survey], and text classification [TC_survey]. However, these models are typically opaque, limiting interpretability, an essential factor for trust, fairness, and compliance [interpretability_1, interpretability_2]. To address the limitations of black-box models, post-hoc explanation methods such as LIME [LIME] and SHAP [SHAP] have been proposed. However, they often lack faithfulness and consistency, offering only approximate and sometimes inaccessible justifications [interpretability_3, interpretability_4], especially for non-technical users. This has led to growing interest in interpretable models that provide transparency by design [interpretability_5].
@@
Prototype-Based Networks (PBNs) follow the paradigm of inherent interpretability, mimicking human reasoning via similarity to class-representative prototypes [cognitive]. While effective in vision [proto_img1, proto_img2] and recently adapted to NLP [10.5555/3648699.3648963-ProtoryNet, wen-etal-2025-gaprotonet], most PBNs are limited to single-label, sentence- or document-level tasks. This limits their use in real-world, multi-label classification scenarios, e.g., “The view was breathtaking, but the staff was rude” expresses opposing sentiments, challenging coarse-grained prototype models [sourati-etal-2024-robust].
@@
Efforts to improve interpretability in text classification often come at the expense of predictive strength. Transformer-based models capture rich semantic structures and typically outperform simpler methods in benchmarks [10.1145/3495162-text-class-survey], yet their lack of transparency limits trust in sensitive applications. Prototype-driven approaches [protolens, wen-etal-2025-gaprotonet] provide clearer, case-based explanations that are easier for users to follow. However, they may sacrifice raw accuracy or struggle to scale to complex domains. This creates a persistent issue: optimizing solely for accuracy risks overlooking fairness and accountability, while prioritizing interpretability may reduce competitiveness in performance-driven evaluations. The issue becomes even more pronounced in multi-label classification [pontiki2016semeval, schouten2015survey], where models must balance fine-grained predictive capability with explanations that remain faithful and comprehensible.
@@
To address these limitations, we propose ProtoSiTex, a semi-interpretable prototype-based framework for fine-grained multi-label classification. ProtoSiTex effectively handles overlapping sentiments and multi-label scenarios via three key innovations: (a) adaptive prototype learning using multi-head attention for subsentence-level semantic alignment, (b) a dual-phase alternate training strategy that combines unsupervised prototype discovery with supervised classification to improve generalization and reduce overfitting, and (c) a hierarchical loss function that enforces consistency across subsentence, sentence, and document levels for structured and faithful predictions. We use the term semi-interpretable to denote models that provide faithful, prototype-based explanations for their predictions while still relying on latent neural mechanisms (e.g., attention, embeddings, hierarchical aggregation) that are not inherently transparent. Our key contributions are:
@@
(i) A semi-interpretable classification framework using adaptive prototypes and multi-head attention: We introduce ProtoSiTex, a prototype-based model primarily for multi-label classification with interpretability. Unlike prior PBNs limited to sentence or document granularity, ProtoSiTex learns adaptive prototypes aligned with subsentence-level class patterns, enabling it to disentangle overlapping or conflicting sentiments. Leveraging multi-head attention, it captures diverse semantic interactions between text segments and prototypes, enhancing both alignment and interpretability.
@@
(ii) Dual-phase alternate training strategy and hierarchical loss for robust learning and semantic consistency: To ensure generalizable and interpretable prototypes, we propose a dual-phase alternate learning strategy: (a) an unsupervised prototype discovery phase to promote diversity and semantic coherence via clustering, and (b) a supervised classification phase that maps prototypes to multiple labels. This decoupling reduces overfitting by allowing periodic prototype refinement. We also introduce a hierarchical loss function to enforce consistency across subsentence, sentence, and document levels, enhancing structured reasoning, predictive accuracy, and explanation fidelity.
@@
(iii) New dataset and extensive evaluation: We curate a hotel reviews dataset (HR) annotated at the subsentence-level with multi-label aspect tags, addressing the limitations of existing sentence or document-label benchmarks. To evaluate ProtoSiTex, we conducted extensive experiments on three datasets: (a) HR (multi-label classification), (b) IMDb (binary classification), and (c) TweetEval (multi-class classification). Across all datasets, ProtoSiTex consistently outperformed existing interpretable baselines and achieved performance on par with state-of-the-art black-box models, showing its effectiveness in terms of accuracy, robustness, and interoperability.
The rest of the paper is organized as follows: Section II briefly reviews the literature, Section III presents our proposed approach, Section IV details the experiments, and Section V concludes with directions for future work.
II Related Work
This section briefly surveys related work and positions our study within the literature.
@@ II Related Work
Interpretable Text Classification: Deep models like CNNs, RNNs, and transformers have achieved strong performance across NLP tasks [10.1145/3495162-text-class-survey], but their opaque nature hinders transparency. Post-hoc explanation methods, such as LIME [LIME] and SHAP [SHAP], aim to interpret predictions by approximating local feature importance. However, these techniques operate independently of model training and often produce inconsistent or non-faithful rationales [cesarini2024explainable-posthoc-limitation]. This has spurred interest in interpretable models that learn to classify and explain simultaneously.
@@
Prototype-Based Networks (PBNs): PBNs classify instances based on similarity to learned prototype vectors, offering human-aligned explanations via exemplar-based reasoning [snell2017prototypical_CV]. Early works such as ProSeNet [10.1145/3292500.3330908-ProsNet] and ClassVector [10.1145/3318299.3318307-ClassVector] introduced prototype learning for document-level, single-label tasks with limited interpretability. Subsequent methods like ProtoryNet [10.5555/3648699.3648963-ProtoryNet] and ProtoLens [protolens] learned finer-grained prototypes at the sentence or subsentence level. GAProtoNet [wen-etal-2025-gaprotonet] integrated graph-attention with prototypes for relational reasoning. Yet, most PBNs focus on single-label classification, lacking the granularity and compositionality required for multi-label classification.
@@
Multi-Label Text Classification (MLTC): MLTC involves assigning multiple labels to a document, making it more challenging than single-label tasks due to inherent label dependencies and data imbalance [multi_label_tpami]. Early approaches relied primarily on problem transformation methods, such as binary relevance, label power set, and classifier chains, as well as algorithm adaptation methods, including multi-label SVMs and KNNs [multi_label_methods]. Models based on CNNs and RNNs improved text representation learning [multi_label_RNN, multi_label_cnn], while GCNs were introduced to explicitly capture label correlations [multi_label_gcn]. More recently, transformer-based architectures, including BERT [devlin-etal-2019-bert] have emerged as strong baselines. Despite these advances, existing methods either treat MLTC as a flat classification problem or rely on complex black-box architectures with limited interpretability. Furthermore, there is hardly any dataset that provides subsentence-level multi-label annotations, limiting empirical benchmarking.
@@
Positioning: ProtoSiTex overcomes these limitations by unifying prototype-based interpretability with fine-grained multi-label classification. Unlike prior works, it learns adaptive prototypes aligned with subsentence semantics and leverages a dual-phase learning strategy, combining unsupervised prototype discovery and supervised hierarchical classification. This design enables structured predictions from subsentence to document level while providing reliable, human-understandable explanations.
III Proposed Methodology
@@
Figure 1: Proposed solution architecture: ProtoSiTex
III-A Problem Formulation
@@
Given a document $\mathcal{D}$ composed of multiple sentences, our goal is to predict its associated class(es), accommodating both single-label and multi-label classification. ProtoSiTex performs document-level prediction while highlighting the specific content segments responsible for each label.
@@
Preprocessing: In our framework, subsentence annotations, though optional, enhance fine-grained interpretability in multi-label tasks. Since manual labeling is costly, we use a few-shot LLM-based annotation engine [ARTICLE_AAAI_2025] to infer them from document-level labels automatically.
@@
Each document $\mathcal{D}$ comprises $n$ sentences, represented as $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots,\mathcal{S}_{n}\}$. Each sentence $\mathcal{S}_{i}$ is further segmented into $t_{i}$ subsentences, defined as $\mathcal{S}_{i}=\{s_{i1},s_{i2},\ldots,s_{it_{i}}\}$, where punctuation marks such as commas and semicolons are used as segmentation cues. The total number of subsentences in a document is denoted by $t=\sum_{i=1}^{n}t_{i}$. To assign fine-grained class labels, we apply the LLM-based engine to generate multi-label annotations at the subsentence level. These annotations are encoded in a binary label matrix $\mathcal{C}\in\{0,1\}^{t\times l}$, where $l$ is the total number of classes. A matrix entry $c_{ij}=1$ indicates that the $i^{\text{th}}$ subsentence is associated with the $j^{\text{th}}$ label. For hierarchical supervision, we aggregate subsentence labels to form sentence-level labels, while document-level labels are directly provided. This bottom-up strategy ensures consistent multi-level annotation. To support structured supervision and hierarchical loss modeling, we construct two binary mapping matrices. The first, $\mathcal{M}^{1}\in\{0,1\}^{1\times n}$, captures the mapping from a document to its sentences. The second, $\mathcal{M}^{2}\in\{0,1\}^{n\times t}$, captures the mapping from each sentence to its associated subsentences. An entry $\mathcal{M}^{2}_{ij}=1$ indicates that sentence or subsentence $j$ belongs to sentence or document $i$, respectively.
III-B ProtoSiTex Architecture
@@
The proposed framework, ProtoSiTex (Fig. 1), consists of three core stages: (i) encoding subsentences into high-dimensional semantic representations using a language model (LM)-based encoder, (ii) employing a dual-phase learning strategy that first discovers representative prototypes through unsupervised learning, then performs supervised classification guided by hierarchical supervision using document-level taxonomy, and (iii) performing inference through prototype-based reasoning applied to the learned embeddings. We detail each component below.
III-B1 Subsentence Representation via LM-based Encoder
@@
Effective text classification relies on context-aware representations. While traditional embeddings (e.g., Word2Vec, GloVe, FastText) are static, LM-based encoders offer dynamic, context-sensitive embeddings but lack interpretability. To balance expressiveness and transparency, we use an LM-based encoder to generate subsentence embeddings.
@@
For each $s_{ij}\in\mathcal{S}_{i}$, the encoder outputs $e_{ij}\in\mathbb{R}^{d}$, producing a feature matrix $\mathcal{E}\in\mathbb{R}^{t\times d}$ over $t$ subsentences. This matrix feeds into the prototype discovery and classification modules, enabling both semantic precision and interpretability. We next describe the dual-phase learning.
III-B2 Dual Phase Learning
The clustering and classification phases aim to learn prototypes and capture multi-granular document representations, respectively, as elaborated below.
Clustering-phase for Prototypes Learning
@@ III-B2 Dual Phase Learning
The first stage of our dual-phase learning aims to discover semantically meaningful and interpretable prototypes that facilitate multi-label classification. To achieve this, we introduce a learnable prototype layer coupled with a multi-head attention (MHA) mechanism, enabling diverse alignment between subsentence embeddings and prototypes.
@@
We initialize $q$ prototypes $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{q}\};\penalty 10000\ \forall p_{i}\in\mathbb{R}^{d}$ using random initialization, which empirically outperforms K-means, Xavier, and Kaiming. The value of $q$ is selected via grid search.
Subsentence embeddings $\mathcal{E}\in\mathbb{R}^{t\times d}$ are projected into the prototype space using MHA, which learns flexible, context-aware alignments. For each attention head $i$, we compute query $\mathcal{Q}_{i}=\mathcal{P}{\cdot}\mathcal{W}^{{Q}_{i}}$, key $\mathcal{K}_{i}=\mathcal{E}{\cdot}\mathcal{W}^{{K}_{i}}$, and value $\mathcal{V}_{i}=\mathcal{E}{\cdot}\mathcal{W}^{{V}_{i}}$ matrices, where $\mathcal{W}^{Q_{i}},\mathcal{W}^{K_{i}},\mathcal{W}^{V_{i}}\in\mathbb{R}^{d\times d}$ are learnable parameters. Scaled dot-product attention computes the alignment:
@@
$\mathcal{H}_{i}=\sigma\left(\left(\mathcal{Q}_{i}{\cdot}\mathcal{K}_{i}^{\top}\right)/\sqrt{d}\right){\cdot}\mathcal{V}_{i};\hskip 16.38895pt\mathcal{H}=\left(\mathcal{H}_{1}\penalty 10000\ \lVert\penalty 10000\ \mathcal{H}_{2}\penalty 10000\ \lVert...\lVert\penalty 10000\ \mathcal{H}_{h}\right){\cdot}\mathcal{W}^{h}$ (1)
@@
where, $\sigma$ denotes softmax function. The resulting matrix $\mathcal{H}\in\mathbb{R}^{q\times d}$ encodes prototype-aware representations of subsentences, which are used in the subsequent classification phase to associate prototypes with class labels.
Classification-phase for Learning Multi-granular Document Details
@@
In the second phase, we train the classifier using prototype embeddings $\mathcal{P}$, subsentence embeddings $\mathcal{E}$, and prototype-aware representations $\mathcal{H}$. To enhance generalization, we apply multi-granular supervision without separate classifiers, enabling two-way explainability: from subsentences to prototypes and prototypes to class-labels. To map prototypes to labels, we pass $\mathcal{H}$ through two dense layers with ReLU followed by a softmax function, producing prediction scores $\tilde{\mathcal{G}}_{1}\in\mathbb{R}^{q\times l}$:
@@
$\footnotesize\tilde{\mathcal{G}}_{1}={\sigma}\left(\mathcal{W}_{3}{\cdot}ReLU\left(\mathcal{W}_{2}{\cdot}ReLU\left(\mathcal{W}_{1}{\cdot}\mathcal{H}\right)\right)\right)$ (2)
@@
Using $\tilde{\mathcal{G}}_{1}$, we derive sentence-level predictions $\tilde{\mathcal{G}}_{2}=f_{clamp}\left(\mathcal{M}^{2}{\cdot}\mathcal{Y}^{\top}{\cdot}\tilde{\mathcal{G}}_{1}\right)$ $\in\mathbb{R}^{n\times l}$ by aggregating subsentence contributions. This is done via the binary prototype-to-subsentence assignment matrix $\mathcal{Y}$, computed by applying cosine similarity between prototypes and subsentences, and sentence-to-subsentence mapping $\mathcal{M}^{2}$.
Finally, document-level predictions $\tilde{\mathcal{G}}_{3}=f_{clamp}\left(\mathcal{M}^{1}{\cdot}\tilde{\mathcal{G}}_{2}\right)\in\mathbb{R}^{1\times l}$ are derived by aggregating sentence-level scores via document-sentence mapping $\mathcal{M}^{1}$.
The $f_{clamp}(\cdot)$ function restricts the elements of a matrix to the $[0,1]$ range by applying element-wise clipping, setting any negative value to 0 and any value above 1 to 1.
@@
We generate dynamic ground-truths for training. Prototype-label associations $\mathcal{G}_{1}$ are computed by first measuring cosine similarity between prototypes and subsentence embeddings to form $\mathcal{X}=\mathcal{P}{\cdot}\mathcal{E}^{\top}/(\lVert\mathcal{P}\rVert\lVert\mathcal{E}^{\top}\rVert)\in\mathbb{R}^{q\times t}$ that is binarized into $\mathcal{Y}$ as:
@@
$\footnotesize\mathcal{Y}_{ij}=\begin{cases}1,&\text{if }i=\arg\max_{k\in\{1\penalty 10000\ ,\dots,\penalty 10000\ q\}}\mathcal{X}_{kj},\\ 0,&\text{otherwise}\end{cases}$ (3)
@@
The binarized $\mathcal{Y}\in\{0,1\}^{q\times t}$ is then multiplied with the subsentence label matrix $\mathcal{C}\in\{0,1\}^{t\times l}$ to yield:
$\footnotesize\mathcal{G}_{1}=\mathcal{Y}\cdot\mathcal{C};\qquad\mathcal{G}_{2}=f_{clamp}(\mathcal{M}^{2}\cdot\mathcal{C})$ (4)